Circular No. 8497

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V382 NORMAE = NOVA NORMAE 2005 W. Liller, Viña del Mar, Chile, reports his discovery of a possible nova

on hypersensitized Technical Pan films taken with an 85-mm f/1.4 Nikon lens and an orange filter. Liller gave the position of the new object as α = 16^{h}19^{m}.8, δ = −51^{o}35′ (equinox 2000.0), and he provided the following red magnitudes: Mar. 9.37 UT, [11; 13.309, 9.4; 18.320, 8.9. Liller additionally obtained the following CCD magnitudes with a 0.20-m Schmidt camera: Mar. 20.326, V = 10.10 ± 0.05, B = 10.90 ± 0.06; Mar. 22.302, V = 10.85 ± 0.03. L. A. G. Monard, Pretoria, S. Africa, reports the following precise position of this object from an unfiltered CCD image taken by himself: α = 16^{h}19^{m}45^{s}.67, δ = −51^{o}36′07′′.2. Magnitudes from Monard: Mar. 20.038 UT, V = 9.8; 20.040, R_{c} = 8.5. N. N. Samus and E. Kazarovets, Institute of Astronomy, Russian Academy of Sciences, inform us that the designation V382 Nor has been given to this nova.

A. Ederoclite, E. Mason, and T. H. Dall, European Southern Observatory (ESO), report spectroscopic confirmation of Liller’s nova with the 3.6-m ESO telescope (+ HARPS; range 380–690 nm; resolution 100000). The nova appears dominated by very strong Hα emission. Low-ionization emissions lines from Fe II (multiplets 27, 28, 37, 38, 42, 48, 49, and 74) and Na I are clearly visible. The Balmer and Fe II lines are flanked by weak P-Cyg profiles that indicate an average expansion velocity of 1100 ± 100 km/s. The Na I interstellar lines appear saturated, thus suggesting a high absorption/reddening.

POSSIBLE SUPERNOVA IN NGC 4656 D. Rich, Hampden, ME, reports his discovery of a possible supernova

(mag ∼ 18.0) on unfiltered CCD frames taken with a 0.31-m reflector on Mar. 21.14 and 22.15 UT. The new object is located at α = 12^{h}43^{m}45^{s}.84, δ = +32^{o}06′15′′.0 (equinox 2000.0), which is ∼ 151′′ west and 230′′ south of the center of NGC 4656. Nothing is visible at this location on CCD frames taken by Rich on 2004 Feb. 27.24 and Mar. 29.25 (limiting mag ∼ 18.7) or on Palomar Sky Survey plates from 1990 (limiting red mag 20.1 and limiting blue mag 19.7). H. Yamaoka, Kyushu University, writes that K. Itagaki, Teppo-cho, Yamagata, Japan, confirmed the new object on his unfiltered archival CCD images taken with a 0.60-m reflector on Mar. 19.751 (red mag 18.5) and 21.678 (at red mag 18.3), and measured the following position end figures from the latter image: 45^{s}.80, 14′′.7. Nothing was detected on an image taken by Itagaki on Mar. 9.612 to red mag 18.5.

2005 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green